Aunt Lucy is pricing cakes for a baby shower she is throwing. She wants one large cake shaped like a duckling and also some cupcakes in pastel colors. Fairview Bakery charges $3 for each cupcake, plus $25 for the large cake. Terrell's Sweet Shoppe charges $57 for the large cake and $1 for each cupcake. If Aunt Lucy orders a certain number of cupcakes, the cost will be the same at either bakery. How many cupcakes would that include? What would the total cost be?
1
Expert's answer
2020-04-07T17:18:44-0400

Let X be the number of cupcakes.

Then 25+3*X=57+X because the cost will be the same at either bakery.

Solving the equation we get

2*X=32,

X=16.

Thus, the number of cupcakes is 16.

$57+$16=$73.

The total cost is $73.